### v1

# 1]Ling Team, Ant Group

Abstract
Recent advances in learning rate (LR) scheduling have demonstrated the effectiveness of decay-free approaches that eliminate the traditional decay phase while maintaining competitive performance. Model merging techniques have emerged as particularly promising solutions in this domain. We present Warmup-Stable and Merge (WSM), a general framework that establishes a formal connection between learning rate decay and model merging. WSM provides a unified theoretical foundation for emulating various decay strategies—including cosine decay, linear decay and inverse square root decay—as principled model averaging schemes, while remaining fully compatible with diverse optimization methods. Through extensive experiments, we identify merge duration—the training window for checkpoint aggregation—as the most critical factor influencing model performance, surpassing the importance of both checkpoint interval and merge quantity. Our framework consistently outperforms the widely-adopted Warmup-Stable-Decay (WSD) approach across multiple benchmarks, achieving significant improvements of +3.5% on MATH, +2.9% on HumanEval, and +5.5% on MMLU-Pro. The performance advantages extend to supervised fine-tuning scenarios, highlighting WSM’s potential for long-term model refinement.
[…]
[3.8cm]assets/gsai.png
1 Introduction
In large language model (LLM) pre-training, learning rate (LR) scheduling plays a pivotal role, critically impacting training stability, convergence speed, and final model performance (Jin et al., 2023; Gotmare et al., 2019). Conventional LR schedules dynamically adjust the LR based on training progress, with the cosine decay schedule emerging as a widely-adopted approach (Kaplan et al., 2020; Hoffmann et al., 2022). This method features an initial warm-up phase followed by cosine-based decay, both constrained by a predetermined total training duration. Consequently, any extension of the training process, such as incorporating new data, necessitates a complete restart from the beginning to recalibrate the entire decay curve.
To address this inflexibility, the Warmup-Stable-Decay (WSD) strategy (Hu et al., 2024) inserts a stable training phase with a constant LR between the warmup and decay phases. This schedule provides two key advantages: (1) flexible initiation of the decay phase independent of total step count, and (2) complete decoupling from fixed training durations. The WSD approach has demonstrated significant effectiveness, evidenced by its adoption in several recent LLMs including DeepSeek-V3 (DeepSeek-AI et al., 2024) and ERNIE 4.5 (ERNIE-Team, 2025). However, WSD introduces new scheduling requirements: researchers must manually decide when to initiate the decay, how many tokens to allocate for it, and which decay function (e.g., cosine, linear) to employ. Furthermore, if training needs to be extended after the decay has commenced, one must roll back the training to the state preceding the decay phase and re-design the decay strategy. This dependency on a manually configured schedule counteracts the goal of a fully autonomous and continuous training process.
Figure 1: Comparison between WSM and mainstream learning rate scheduling strategies.
To minimize scheduling complexity, recent research has investigated alternative approaches that completely eliminate the decay phase from LR schedules (Defazio et al., 2024; Song et al., 2025; Zhang et al., 2025). A prominent direction in this field explores weight averaging (a.k.a., model merging) techniques. Empirical results demonstrate that simply maintaining a constant LR combined with standard weight averaging strategies—such as exponentially weighted averaging (EWA)—can achieve performance competitive with WSD-based schedules (Li et al., 2025). Building upon this line of research, we make three key extensions:
(1) We present Warmup-Stable and Merge (WSM), a simple yet general framework for LR scheduling. By formalizing the connection between LR decay and checkpoint merging, we demonstrate that WSM can be instantiated to emulate various decay strategies—including cosine decay, linear decay and inverse square root (1-sqrt) decay. Our framework provides a principled approach to convert any LR decay method into a theoretically approximate model averaging implementation. This contrasts with prior work (Defazio et al., 2024; Song et al., 2025; Li et al., 2025), which has largely focused on analysis of specific averaging strategies and their optimization properties. Notably, our framework is optimizer-agnostic, enabling seamless integration with various optimization algorithms (e.g., SGD, Adam) without requiring modifications on the underlying training pipeline.
(2) Through extensive experiments, we systematically investigate key factors in instantiating this framework, including merge methods, merge frequency, duration, granularity, and the compatibility of merging and decay strategies. Our findings reveal that merge duration—the training period covered by the merged checkpoints—emerges as the most critical factor influencing model performance, with significantly greater impact than both the checkpoint interval and the number of merged models.
(3) The proposed decay-free LR schedule delivers substantial performance gains. Extensive empirical evaluation shows consistent improvements across multiple benchmarks: +3.5% on MATH, +2.9% on HumanEval, and +5.5% on MMLU-Pro—a significant advance over the WSD method, while prior work often only matches WSD’s performance. Moreover, these benefits naturally extend to the post-training stage, demonstrating our method’s potential for sustained improvements in long-term model refinement.
The proposed WSM framework presents a promising direction for developing effective decay-free LR schedules. Our results demonstrate consistent performance advantages across both pre-training and fine-tuning stages. Notably, WSM enables the implementation of sophisticated decay-like methods through a simple and stable optimization approach, combining the benefits of strategic scheduling with robust training dynamics.
2 Preliminary
Typically, existing mainstream LR schedulers consist of two phases: an initial warm-up followed by decay. During the warm-up, the LR increases linearly from a small value to a peak value, $lr_{peak}$, over $T_{warmup}$ steps. This helps stabilize the optimization process in the early stages of training. Following the warm-up, the LR gradually decreases according to a predefined function, such as cosine, linear, or inverse square root decay. These schedules share a critical limitation: they require the total number of training tokens (or steps), $T_{max}$, to be known in advance. For instance, the prevalent cosine LR schedule is formulated as:
$lr(t)=\begin{cases}lr_{peak}\cdot\frac{t}{T_{warmup}}&\text{if }t<T_{warmup}\\ \frac{1}{2}lr_{peak}\left(1+\cos\left(\frac{\pi(t-T_{warmup})}{T_{max}-T_{warmup}}\right)\right)&\text{if }t\geq T_{warmup}\end{cases}$
The Warmup-Stable-Decay (WSD) LR schedule introduces a stable phase between warm-up and decay, maintaining a constant LR at its peak ($lr_{peak}$), which is defined as:
$lr(t)=\begin{cases}lr_{peak}\cdot\frac{t}{T_{warmup}}&\text{if }t<T_{warmup}\\ lr_{peak}&\text{if }T_{warmup}\leq t<T_{decay\_start}\\ \operatorname{decay\_function}(t)&\text{if }T_{decay\_start}\leq t\leq T_{max}\end{cases}$
As shown, WSD eliminates the need for manual LR tuning during the stable phase while supporting multiple decay attempts from its endpoint—without requiring a reset to the initial state. Despite this flexibility, it still requires predefined decay-phase settings—such as the decay start step ($T_{decay\_start}$), decay function ($\operatorname{decay\_function}$), and total training steps ($T_{max}$).
3 The Proposed Methodology
In this section, we first establish the theoretical connection between checkpoint merging and LR decay, formalize our proposed WSM (Warmup-Stable and Merge) schedule, and compare it with the widely used WSD schedule.
3.1 Theoretical Connection Between LR Decay and Checkpoint Merging
The core idea of checkpoint merging in this work is to take an ordered list of checkpoints, $[\theta_{n},\theta_{n+1},\dots,\theta_{n+k}]$, and apply a merge function to generate a single model $\hat{\theta}_{n+k}$. Here, $\theta_{i}\in\mathbb{R}^{d}$ represents the model’s parameter vector at the $i$-th training iteration. The most general form is a weighted average of the checkpoints:
$\hat{\theta}_{n+k}=\sum_{j=0}^{k}c_{j}\theta_{n+j}$ (1)
where $\{c_{j}\}$ are non-negative weights that sum to one, i.e., $\sum_{j=0}^{k}c_{j}=1$.
This formulation obscures a deeper connection to the training dynamics. We can reveal this connection by expressing each checkpoint in terms of an initial checkpoint $\theta_{n}$ and the subsequent gradient updates. For simplicity, we assume the updates between checkpoints at different time steps are independent and ignore optimizer states. Let $g_{i}$ be the gradient update vector (including the LR) at step $i$, such that the model updates as $\theta_{i+1}=\theta_{i}-g_{i}$. Thus, an intermediate checkpoint $\theta_{n+j}$ can be expressed as the sum of an initial state $\theta_{n}$ and the sequence of negative gradient updates that followed:
$\theta_{n+j}=\theta_{n}-\sum_{l=1}^{j}g_{n+l-1}$ (2)
Figure 2: Visualization of checkpoint merging weight distributions and their corresponding decay functions over a merge duration of 10 checkpoints (larger checkpoint index denotes more recent checkpoints). (a) Exponential Moving Average (EMA) weights, exhibiting convex decay characteristics; (b) Uniform averaging weights, demonstrating linear decay behavior; (c) and (d) Our theoretically derived weights, designed to approximate cosine and 1-sqrt decay patterns. Note: We draw the decay curves smoothly for better illustration, although they are actually step-like due to their discrete nature.
Furthermore, we can substitute Eq. 2 into the general merging formula (Eq. 1) and then rearrange the double summation by changing its order. A gradient update $g_{n+i-1}$ is included in the sum for all involved checkpoints $\theta_{n+j}$ where $j\geq i$.
$\hat{\theta}_{n+k}$ $=\sum_{j=0}^{k}c_{j}\left(\theta_{n}-\sum_{l=1}^{j}g_{n+l-1}\right)=\theta_{n}-\sum_{i=1}^{k}\left(\sum_{j=i}^{k}c_{j}\right)g_{n+i-1}$ (3)
This shows that a weight coefficient of $\sum_{j=i}^{k}c_{j}$ is applied to the gradient update $g_{n+i-1}$ from step $n+i-1$. By defining a new set of weights for the gradient updates, $w_{i}=\sum_{j=i}^{k}c_{j}$, we arrive at the final equivalent expression for checkpoint merging:
$\hat{\theta}_{n+k}=\theta_{n}-\sum_{i=1}^{k}w_{i}\cdot g_{n+i-1}$ (4)
This equation shows that merging checkpoints with weights ${c_{j}}$ is equivalent to applying a synthetic decay schedule defined by weights ${w_{i}}$ to the gradients accumulated after the base checkpoint $\theta_{n}$, where a mapping exists between the merge weights ${c_{j}}$ and the effective learning rates ${w_{i}}$. We therefore propose Theorem 3.1, which can approximate monotonically decreasing decay curve. Figure 2 illustrates various checkpoint merging weights and their corresponding decay curves. Additional proof details are provided in Appendix B.
Theorem 3.1 (Checkpoint Weight Derivation from Gradient Decay Schedule).
Given a desired sequence of gradient decay coefficients $\{w_{i}\}_{i=1}^{k}$ that is monotonically non-increasing and bounded, $1\geq w_{1}\geq w_{2}\geq\dots\geq w_{k}\geq 0$, the corresponding non-negative checkpoint weights $\{c_{j}\}_{j=0}^{k}$ required to satisfy Eq. 4 are uniquely determined by:
$\begin{cases}c_{k}=w_{k}\\ c_{j}=w_{j}-w_{j+1},&\text{for }j\in[1,k-1]\\ c_{0}=1-\sum_{j=1}^{k}c_{j}=1-w_{1}\end{cases}$ (5)
3.2 Emulating LR Decay through Checkpoint Merging
The central hypothesis of WSM is that the optimization benefits of LR decay can be decoupled from the live training process and instead be effectively achieved through the merging of model checkpoints (Kingma and Ba, 2015; DeepSeek-AI et al., 2024; Li et al., 2025). The WSM simplifies the LR schedule by completely eliminating the decay phase:
$lr(t)=\begin{cases}lr_{peak}\cdot\frac{t}{T_{warmup}}&\text{if }t<T_{warmup}\\ lr_{peak}&\text{if }t\geq T_{warmup}\\ \end{cases}$
The WSM pipeline, detailed in Algorithm 1, operates in two primary phases. It begins with a standard warmup phase, where the learning rate linearly increases to its peak value, $lr_{peak}$. Following this, the process enters the main stable training phase, where the learning rate is held constant. After a specified step $T_{switch}$, the model can transition from the general pre-training data $D$ to a smaller, high-quality annealing dataset $D_{anneal}$. This allows the “annealing” to be focused on a curated data mixture, aligning with practical pre-training methodologies. Throughout this stable phase, checkpoints are periodically saved. Concurrently, an asynchronous merging process continuously fetches the last $n$ checkpoints from storage and combines them into model $W_{merged}$. Specifically, for the $\operatorname{Merge}(\cdot)$ operation, we can select various decay strategies to emulate (e.g., the decay curve shape and minimum LR), calculate the corresponding gradient decay coefficients $\{w_{i}\}$, and then derive the checkpoint merging weights $\{c_{i}\}$ based on Theorem 3.1. This merged checkpoint, which emulates the effect of a decay schedule, provides a robust, annealed model without ever altering the live learning rate.
Algorithm 1 The WSM LRS Pre-training Pipeline
Input: Initial model weights $W_{0}$, pre-training data $D$, high-quality annealing data $D_{anneal}$ (optional), peak LR $lr_{peak}$, warmup steps $T_{warmup}$, switch step $T_{switch}$, checkpointing interval $T_{cpt}$, merging window size $n$
Initialize:
$\triangleright$ The merged checkpoint.
$W_{merged}\leftarrow\text{null}$
$\triangleright$ Storage for the base (un-merged) checkpoints.
$\mathcal{C}_{storage}\leftarrow[]$
$\triangleright$ Phase 1: Warmup on Pre-training Data
for $t=1$ to $T_{warmup}$ do
    $lr(t)\leftarrow lr_{peak}\cdot(t/T_{warmup})$
    $W_{t}\leftarrow\text{Update}(W_{t-1},D,lr(t))$
end for
$\triangleright$ Phase 2: Stable Training and Merging
for $t=T_{warmup}+1$ to … do
    $lr(t)\leftarrow lr_{peak}$
    $\triangleright$ Select dataset for the current step
    if $t>T_{switch}$ and $D_{anneal}$ is available then
         $D_{current}\leftarrow D_{anneal}$
    else
         $D_{current}\leftarrow D$
    end if
    $W_{t}\leftarrow\text{Update}(W_{t-1},D_{current},lr(t))$
    if $t\pmod{T_{cpt}}==0$ then
         $\triangleright$ Save the checkpoint from the main training process.
[…]
         $\triangleright$ Asynchronously Update the Merged checkpoint
         $C_{latest}\leftarrow\text{GetLastNCheckpoints}(\mathcal{C}_{storage},n)$
         if $\text{len}(C_{latest})==n$ then
             $\triangleright$ Update the merged checkpoint for evaluation.
             $W_{merged}\leftarrow\text{Merge}(C_{latest})$
         end if
    end if
end for
return The stored base checkpoints in $\mathcal{C}_{storage}$ and the merged checkpoint $W_{merged}$.
3.3 Discussion
Choice between offline and online merging
Initially, we employ an offline, checkpoint-based approach as a powerful exploration framework to discover optimal annealing strategies without impacting the primary training run. By preserving a history of all discrete checkpoints, practitioners can systematically evaluate how different factors influence the final model, including the choice of merging algorithm (simulating various decay curves like linear or cosine) and the annealing duration (by varying the merge window size, $n$). This multi-faceted exploration is impossible with a standard online method, such as an Exponential Moving Average (EMA), which hard-codes a single annealing path into the training process. Crucially, once this offline exploration identifies a superior, fixed strategy (e.g., “an average of the last 4 checkpoints”), it can then be operationalized as a simple and efficient online process using a sliding window.
Practicality and complexity
Compared to discovering an optimal decay schedule through multiple, resource-intensive training runs, the merging operations in the WSM strategy save substantial computational resources. The main complexity introduced is storage overhead. For the offline merging approach, which is ideal for exploration, one must store a history of checkpoints, with the total number being the total training tokens divided by the checkpointing interval. However, this is a manageable trade-off, as this storage footprint represents a minor fraction of a typical pre-training budget. Furthermore, once a superior strategy is identified, or in scenarios with extreme storage constraints, an online merging approach can be used. This method minimizes the storage footprint by maintaining only a small, fixed-size window of $n$ checkpoints. Our experiments confirm that a relatively small window (e.g., $n=12$) is sufficient to achieve strong results.
Comparison with Li et al. (2025).
We notice a concurrent work (Li et al., 2025) studied model merging in pre-training. We highlight the differences between our work and theirs as follows: (1) Motivation: We frame checkpoint merging as a novel LR scheduling mechanism, with the primary goal of discovering a schedule that outperforms strong baselines like WSD. In contrast, their work treats model merging as a standalone pre-training technique, evaluating its value by comparing the performance of the merged model against its before-merge checkpoints. (2) Methodology: Our approach is theory-driven. We first establish a formal theoretical connection between LR decay and checkpoint merging (Section 3.1) and then leverage merge weights to simulate various decay curves. They mainly conducts empirical studies of merging algorithms, proposing heuristic methods including WMA, SMA, and EMA. (3) Key Findings: Through extensive experiments, we identify the merge duration as the most critical hyperparameter influencing final model performance. Ultimately, our method surpasses the performance of WSD-annealed baseline (Section 4.2). In contrast, their work concludes that checkpoint merging can effectively match the performance of an annealed model. The findings from our work and that of Li et al. (2025) are complementary, collectively providing practical insights into checkpoint merging for LLM pre-training.
4 Experiment
4.1 Experiment Setup
The model we used for the experiment is Ling-mini, a standard MoE model with a total of 16.3 billion parameters and 1.4 billion active parameters. Ling-mini adopts a fine-grained expert configuration, consisting of 1 dense layer and 19 MoE layers. Each MoE layer includes 256 experts, with 8 experts being activated for each token, along with one additional shared expert. We utilized the AdamW optimizer (Loshchilov and Hutter, 2019), and the hyperparameters are set to $\beta_{1}=0.9$ and $\beta_{2}=0.95$, with 0.1 weight decay applied. Through preliminary scaling laws experiments, we set the peak LR and batch size to 4.78e-4 and 2048, respectively.
We begin with a checkpoint pretrained on 10.2 trillion tokens total - comprising 10 trillion tokens at standard 4,096 length sequences plus 200 billion long-context tokens - all trained using a stable, constant LR. Then, we introduce a specialized high-quality annealing dataset and branch the training into two distinct strategies for an additional 400B tokens to compare their effectiveness: (1) We apply a conventional LR decay schedule to the model. This branch serves as our baseline, representing the standard Warmup-Stable-Decay (WSD) methodology. (2) We continue training with the same constant LR. The final model is then produced by merging the checkpoints saved during this stable phase. This branch represents our proposed WSM schedule. Unless otherwise specified, we save a checkpoint every 25B tokens and use mean averaging to merge the most recent checkpoints. Comprehensive details regarding our model architecture, specific training parameters, dataset composition, and evaluation protocols are provided in Appendix A.
4.2 Overall Performance of WSM Schedule
Figure 3: Comprehensive performance comparison (overall and by category) between our WSM schedule (via checkpoint merging) and standard WSD scheduling (via learning rate decay). Both approaches are initialized from the same pretrained checkpoint (10.2T tokens with constant LR). Notably, while WSD requires predetermined decay strategy (e.g., decay over 400B tokens in this study), WSM eliminates such constraints, enabling seamless training continuation (gray regions) and flexible decay behavior approximation.
Immediate effects during pre-training
In this section, we present a comprehensive performance comparison between our proposed WSM schedule and the baseline WSD schedule. We evaluate three series of checkpoints: (1) those obtained using the standard LR decay schedule in WSD, (2) checkpoints from the last stable phase of WSM before merging, and (3) our final merged checkpoints from the WSM method using various merge durations (window sizes). The category-wise average results are summarized in Figure 3 and Table 2, with scores for each benchmark provided in Appendix E.
Our first and most significant finding is that the WSM method consistently outperforms WSD across the majority of tasks considered. On average, WSM achieves performance improvements across all benchmark categories. Notably, when comparing the best-performing checkpoint from each strategy, WSM improves upon WSD by an average of 1.3 points. We observe remarkable improvements of up to 2.7 points on MATH, 2.4 points on HumanEval, 2.1 points on MMLU-Pro. These results provide compelling evidence that replacing the LR decay phase used in WSD with the checkpoint merging strategy of WSM is not only a feasible alternative but also a more effective approach for enhancing the diverse capabilities of the final pretrained model.
Long-term implications for post-training
While WSM demonstrates promising results in the pre-training phase, we further investigate its long-term impact on subsequent post-training stages. We apply supervised fine-tuning on checkpoints generated by the WSM and WSD under identical settings for 5 epochs. Results in Table 2 show that this advantage persists beyond the post-training phase.
Table 1: Base model performance comparison. Results are reported based on the checkpoint with the highest average benchmark score.
General Knowledge Language Modeling Math Code Professional Knowledge Overall Average
Base Model WSD 69.06 67.78 57.49 64.88 53.46 62.67
WSM 70.22 68.67 58.81 65.58 56.04 63.95
Improv. +1.68% +1.31% +2.30% +1.08% +4.83% +2.04%
Table 2: Instruct model performance comparison. Results are reported based on the epoch with the highest average benchmark score.
Language Knowledge Math Code Reason Agent Overall Average
Inst Model WSD 81.12 60.00 61.43 58.23 63.21 68.16 62.90
WSM 84.78 61.73 62.28 57.95 64.94 69.33 64.07
Improv. +4.51% +2.88% +1.38% -0.48% +2.74% +1.72% +1.86%
4.3 Empirical Analysis of WSM Schedule
In this section, we conduct a comprehensive empirical study to dissect the WSM schedule. We aim to understand the key factors influencing its performance, its robustness across different training scenarios, its interaction with conventional decay mechanisms, and its broader implications on model dynamics.
4.3.1 Robustness Across Pre-training Process
Beyond applying WSM as a final step on a high-quality dataset, we also evaluated its utility and robustness throughout the entire pre-training lifecycle. To achieve this, we conducted a comparative analysis at various intermediate stages of a long training run, comparing the performance of a model produced by our computationally-frugal WSM merging against one produced by initiating a full, resource-intensive learning rate (LR) decay. As shown in Figure 5(a), although the performance gains over WSD are not as significant as when switching to high-quality data, we find that the performance of models generated via WSM merging consistently and closely mirrors the results of a true LR anneal. In the figure, the gray line represents the base model trained with a constant LR. The blue lines show the performance of WSM models, which were created by mean-merging four checkpoints from within the preceding 100B-token merge duration. The red lines represent full 100B-token decay runs initiated at the 2T, 4T, 6T, 8T, and 10T token milestones. This result establishes WSM as a reliable, high-fidelity proxy for estimating a model’s post-anneal potential at any point during training. Consequently, it can provide effective assessment throughout the pre-training phase, eliminating the need to launch multiple, expensive decay runs to gauge the model’s true strength.
4.3.2 Impact of Merging Algorithm
As derived in Section 3.1, the checkpoint merging process can be viewed as an approximation of a LR decay schedule, where the weighting scheme of the merge is analogous to the functional form of the decay curve. For instance, a simple mean average is analogous to a linear decay curve. An Exponential Moving Average (EMA) would correspond to a convex exponential decay curve. Our prior experiments with the WSD schedule revealed a performance hierarchy among decay curves: concave schedules (e.g., inverse square root) and linear schedules outperform convex schedules (details are provided in Appendix C). Building on these findings, we hypothesize that a merging algorithm designed to approximate decay curves that have been proven effective in WSD scheduling will similarly yield better results.
Based on Theorem 3.1, we experimentally compare three merging algorithms: one using our theorem-generated weights to approximate 1-sqrt decay, another using simple mean averaging, and a third using EMA. Our experimental results in Table 3 validate this hypothesis. While the merge method outperforms decay, the 1-sqrt merge approach shows slight advantages over Mean, and both are markedly better than EMA. This finding reinforces the theoretical connection between checkpoint merging and LR decay, suggesting that the benefits of superior decay schedules can be effectively captured through carefully designed merging weights.
4.3.3 Impact of Merging Duration and Granularity
Figure 4: Merge duration analysis across algorithms.“Merge 4” indicates merging the most recent 4 checkpoints, with each checkpoint saved at 25B-token intervals.
We further investigate in detail the impact of different merge durations (window sizes) on various algorithms. First, different merge durations essentially correspond to decaying with varying amounts of data. As shown in Figure 4, when comparing the best-performing checkpoints across the entire merging trajectory for both mean and 1-sqrt merging algorithms, larger merging windows tend to yield better results. However, this advantage gradually diminishes as the window size increases. This observation aligns with previous practical LR decay experiments, where simply increasing the amount of annealing data shows diminishing returns and may eventually fail to provide further improvement. These findings further strengthen the connection between LR decay and checkpoint merging. For EMA merging, its performance is significantly inferior to other algorithms and shows no clear variation with merge durations. This may indicate that EMA is not an effective merging algorithm (also suggesting that the convex nature may not represent an optimal decay curve).
Then we further investigate the granularity of checkpoint merging, i.e., the interval between saved checkpoints used for merging. Finer-grained merging represents a more precise approximation of the true decay curve. The results in Table 4 show that finer-grained merging tends to achieve better performance. However, frequent checkpoint saving imposes storage overhead, requiring careful trade-off considerations.
Table 3: Impact of merging algorithm.
General Knowledge Language Modeling Math Code Professional Knowledge Overall Average
Decay 1-sqrt 69.06 67.78 57.49 64.88 53.46 62.67
Merging EMA 69.05 67.64 58.81 64.19 54.44 63.01
Mean 70.22 68.67 58.81 65.58 56.04 63.95
1-sqrt 70.27 68.26 59.65 65.70 55.42 64.06
Table 4: Comparison of different saving/merging intervals within an 80B-token merge duration. For example, (5B,16) indicates that saving every 5B tokens while merging the latest 16 checkpoints.
Merging Granularity General Knowledge Language Modeling Math Code Professional Knowledge Overall Average
(5B,16) 69.46 79.95 57.07 64.68 53.82 63.63
(10B,8) 69.23 80.00 57.94 64.39 53.78 63.78
(20B,4) 69.29 80.29 56.79 63.87 54.19 63.36
(40B,2) 68.47 79.83 56.57 63.59 52.20 62.77
(80B,1) 67.47 64.98 55.07 61.69 51.61 60.33
4.3.4 On the Compatibility of Merge and Decay
Given that checkpoint merging effectively simulates LR decay, a natural question arises: can merging and decay be combined to achieve synergistic performance gains? We investigate this by testing two hybrid approaches. (1) Decay-then-Merge: We first apply a standard decay schedule and then merge checkpoints selected from within the decay phase. (2) Merge-then-Decay: We further apply a decay schedule to the resulting merged model.
Figure 5: (a) A comparison of checkpoint merging (WSM) and true LR decay (WSD) across a long pre-training run. (b) Investigating the effect of applying merging within a decay phase (Decay-then-Merge). (c) Investigating the effect of applying LR decay after checkpoint merging (Merge-then-Decay).
As shown in Figures 5(b) and 5(c), the hybrid approach failed to yield any improvement in either configuration, although the Merge-then-Decay model showed better performance at the beginning of its training. For the Decay-then-Merge experiment (Figure 5(b)), the blue stars represent the results of merging checkpoints selected along the decay trajectories (red lines), which were initiated at various pre-training milestones (4T, 6T, 8T, 10T). For the Merge-then-Decay experiment (Figure 5(c)), we compare a decay run initiated from a WSM model (blue line)—created by merging four checkpoints from the 9.8T to 10T token interval—against a standard decay baseline initiated from a single 10T checkpoint (red line). These results suggest that checkpoint merging and LR decay are not complementary but rather alternative pathways to a similar optimization objective.
4.3.5 Impact on MoE Load Balancing
Table 5: Impact on MoE load balancing. The WSM strategy demonstrates improved expert utilization (lower load balancing violation scores) with a slightly higher language modeling loss.
language modeling loss mean_global_max_violation mean_global_min_violation
WSD 0.675 0.601 0.322
WSM 0.697 0.545 0.201
[…]
4.4 Parameter Trajectories and Model Performance: Constant LR with Merge vs. LR Decay
To visually analyze the correlation between model parameters and performance during training, we employ t-SNE (van der Maaten and Hinton, 2008) dimensionality reduction to project the weight matrix of a specific layer into the 2-dimensional embedding space. This projection is combined with performance evaluations to generate the composite contour map illustrated in Figure 6. The directional arrows in the visualization explicitly illustrate the parameter trajectories across both constant and decay phases. Our experimental analysis revealed two principal findings: 1) During the decay phase, model parameters gradually converge toward the merged model solution space, which achieves superior performance compared to checkpoints with a constant LR. 2) The LR reduction in the decay phase enables more precise parameter refinement than the expansive exploration observed under constant LR. This controlled convergence facilitates localization of nearby optimal solutions. We confirm the speculation of previous work Wen et al. (2024), and formalize this dynamic with an analogy: constant LR training resembles traversing a “canyon” with oscillating steps, while merging resembles finding a “river” at the canyon floor that guides efficient convergence.
Figure 6: Visualization of Model Parameters and Performance Contour Lines. Points connected by directional arrows represent parameter trajectories during the constant and decay phases, respectively. The red star indicates the solution space of the merged model.
5 Related Work
5.1 Learning Rate Schedule
Learning rate (LR) scheduling is critical for training performant models (Jin et al., 2023; Gotmare et al., 2019). Classic schedules like Cosine (Kaplan et al., 2020; Ibrahim et al., 2024) or Linear (Defazio et al., 2023) decay adjust the LR based on a predefined total training duration, which is inflexible for continual training. The Warmup-Stable-Decay (WSD) schedule (Hu et al., 2024) addresses this by introducing a stable LR phase after warmup, decoupling the eventual decay from a fixed training length and offering greater flexibility for long or continuous training runs. More recently, researches have explored “schedule-free” methods to eliminate the decay phase entirely, which maintain a constant LR and instead leverage weight averaging techniques (Defazio et al., 2024; Song et al., 2025; Zhang et al., 2025). Builds upon these schedule-free principles, our work propose to replace WSD’s decay phase with a post-hoc checkpoint merging operation instead of specific online averaging strategies. This simplifies the training pipeline and, by formalizing the connection between LR decay and checkpoint merging, allows decay strategy to be theoretically approximated, leading to free offline exploration and enhanced model performance.
5.2 Model Merging
Model merging (Izmailov et al., 2018; Wortsman et al., 2022) has emerged as an efficient paradigm for model construction. This approach achieves effective knowledge transfer and performance improvement through parameter-level integration of multiple models. Model merging is primarily utilized in two distinct scenarios: (1) The integration of knowledge and capabilities from multiple independently trained models into a single parameter set, with the objective of preserving maximal performance from each specialized model (Aakanksha et al., 2025; Ramé et al., 2024); and (2) the merging of checkpoints along a single training trajectory, which functions as a form of Polyak averaging (Polyak and Juditsky, 1992). In this application, the merging operation acts as a smoothing mechanism that reduces noise inherent in stochastic gradient-based optimization (Sanyal et al., 2023; Liu et al., 2024a; Kaddour, 2022; Li et al., 2023; Sandler et al., 2023). While Li et al. (2025) show model merging can achieve performance competitive with decay-based schedule, these techniques have also demonstrated practical utility in industrial-scale LLM development (Grattafiori et al., 2024; DeepSeek-AI et al., 2024; Aakanksha et al., 2025). Our work falls into the second category of merging checkpoints along a single trajectory. We establish a theoretical connection between this operation and learning rate decay and provide a principled approach to convert various LR decay strategies into a theoretically approximate model averaging implementation.
6 Conclusion
In this paper, we have presented WSM, a decay-free LR scheduling approach for LLM pre-training. Our method bridges LR decay and checkpoint merging by establishing the theoretical connection. By eliminating the conventional decay phase, WSM simplifies LR scheduling while reformulating various decay strategies as principled model averaging schemes. Through systematic analysis, we identified merge duration as the most critical factor influencing model performance—outweighing other implementation choices. Extensive experiments have demonstrated WSM’s superiority over traditional WSD schedules, with consistent improvements across multiple benchmarks and sustained benefits during fine-tuning. Due to its optimizer-agnostic design, WSM offers broad applicability without modifying the underlying training pipeline.
As future work, we plan to expand the WSM framework by incorporating additional instantiations of decay strategies, as it can directly convert any decay strategy into a theoretically approximate checkpoint merging scheme. Given its flexibility, we also aim to adapt WSM to more complex tuning scenarios, such as dataset mixture optimization.
Appendix A Experimental Settings
Model architecture
The core architectures of our experimental model are detailed in Table 6. The model is configured with 20 layers and a hidden dimension size of 2048. Except for the first layer, all FFNs layers are replaced with MoE layers. We adopt the GQA attention mechanism (Ainslie et al., 2023) and integrate Rotary Position Embedding (RoPE) (Su et al., 2024), enabling the model to support sequence lengths up to 8K tokens. For parameter initialization, all learnable parameters are randomly initialized using a standard deviation of 0.006. Under this configuration, the model consists of a total of 16.3 billion parameters, of which approximately 1.43 billion are activated for each token during inference.
Table 6: Detailed model architectures.
$n_{layers}$ $d_{model}$ $d_{ffn}$ $d_{expert}$ $n_{heads}$ $n_{kv\_head}$ $E$ $E_{a}$ $E_{s}$ $N$ $N_{a}$
20 2048 5120 512 16 4 256 8 1 16.3B 1.43B
Training hyperparameters
We use the AdamW optimizer (Loshchilov and Hutter, 2019) with hyperparameters set as follows: $\beta_{1}=0.9$, $\beta_{2}=0.95$, and weight decay of 0.1. Gradient clipping (Zhang et al., 2020) norm is set to 1.0. According to the scaling laws for MoE optimal hyper-parameters, the maximum learning rates were set to $3.74e{-4}$. The batch size is set to 2048, and with a maximum sequence length of 8K, each training batch contains 16M tokens.
Pre-training dataset
The training data is sourced from a large-scale multilingual corpus created by the Ling Team, primarily covering English and Chinese, while also including various other languages. This corpus encompasses web text, mathematical materials, programming scripts, published literature, and diverse textual content. To validate model performance, we extracted a 10T-token subset from this corpus for training.
Evaluation setup
To evaluate performance, we consider a diverse suite of downstream tasks designed to provide a holistic assessment of model capabilities. For base model, tasks are grouped into several categories, such as: (a) General Knowledge/Reasoning (e.g., ARC (Bhakthavatsalam et al., 2021), AGIEval (Zhong et al., 2024), OpenBookQA (Mihaylov et al., 2018), BBH (Suzgun et al., 2023), WorldSense (Hong et al., 2025), PIQA (Bisk et al., 2020), hellaswag (Zellers et al., 2019) and KOR-Bench (Ma et al., 2025)) (b) Language Understanding (e.g., race (Lai et al., 2017), SQuAD 2.0 (Rajpurkar et al., 2018), TriviaQA (Joshi et al., 2017), NQ (Kwiatkowski et al., 2019) and winogrande (Sakaguchi et al., 2021)) (c) Professional Knowledge (e.g., MMLU (Hendrycks et al., 2021a), CMMLU (Li et al., 2024a), C-Eval (Huang et al., 2023), MMLU-Pro (Wang et al., 2024), GPQA (Rein et al., 2023) and SuperGPQA (Team et al., 2025)) (d) Math (e.g., GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021b), gaokao (Zhang et al., 2023), GSM-Plus (Li et al., 2024b), mgsm-zh (Shi et al., 2023), CMATH (Wei et al., 2023), MathBench (Liu et al., 2024b), minerva_math (Hendrycks et al., 2021b), college_math (Tang et al., 2024) and cn_middle_school_24) (e) Code (e.g., HumanEval (Chen et al., 2021), LiveCodeBench (Jain et al., 2025), MBPP (Tao et al., 2024), HumanEval_plus (Liu et al., 2023), MBPP_plus (Liu et al., 2023), HumanEval_cn (Peng et al., 2024), HumanEval_fim (Bavarian et al., 2022) and CruxEval (Gu et al., 2024)). For SFT model, the categories and tasks are shown in Table 8 and 9.
Appendix B Details of Theoretical Connection Between Decay and Merging
The core idea of checkpoint merging in this work is to take an ordered list of checkpoints, $[\theta_{n},\theta_{n+1},\dots,\theta_{n+k}]$, and apply a merge function to generate a single model $\hat{\theta}_{n+k}$. Here, $\theta_{i}\in\mathbb{R}^{d}$ represents the model’s parameter vector at the $i$-th training iteration.
The most general form is a weighted average of the checkpoints:
$\hat{\theta}_{n+k}=\sum_{j=0}^{k}c_{j}\theta_{n+j}$ (6)
where $\{c_{j}\}$ are non-negative weights that sum to one, i.e., $\sum_{j=0}^{k}c_{j}=1$.
While intuitive, this formulation obscures a deeper connection to the training dynamics. We can reveal this link by expressing each checkpoint in terms of an initial checkpoint $\theta_{n}$ and the subsequent gradient updates.
We assume that the updates between checkpoints at different time steps are independent. Let $g_{i}$ be the gradient update vector (including the learning rate) at step $i$, such that the model is updated via $\theta_{i+1}=\theta_{i}-g_{i}$. Any checkpoint $\theta_{n+j}$ can therefore be written as the sum of an initial state $\theta_{n}$ and the sequence of negative gradient updates that followed:
$\theta_{n+j}=\theta_{n}-\sum_{l=1}^{j}g_{n+l-1}$ (7)
Substituting Eq. 7 into the general merging formula (Eq. 6) yields:
$\hat{\theta}_{n+k}$ $=\sum_{j=0}^{k}c_{j}\left(\theta_{n}-\sum_{l=1}^{j}g_{n+l-1}\right)$ (8)
$=\left(\sum_{j=0}^{k}c_{j}\right)\theta_{n}-\sum_{j=0}^{k}c_{j}\sum_{l=1}^{j}g_{n+l-1}$ (9)
$=\theta_{n}-\sum_{j=1}^{k}c_{j}\sum_{l=1}^{j}g_{n+l-1}$ (10)
The double summation in Eq. 10 can be rearranged by changing the order of summation. A gradient update $g_{n+i-1}$ is included in the sum for all checkpoints $\theta_{n+j}$ where $j\geq i$.
$\sum_{j=1}^{k}c_{j}\sum_{l=1}^{j}g_{n+l-1}$ $=\sum_{i=1}^{k}\left(\sum_{j=i}^{k}c_{j}\right)g_{n+i-1}$ (11)
This shows that a weight coefficient of $\sum_{j=i}^{k}c_{j}$ is applied to the gradient update $g_{n+i-1}$ from step $n+i-1$. By defining a new set of weights for the gradient updates, $w_{i}=\sum_{j=i}^{k}c_{j}$, we arrive at the final equivalent expression for checkpoint merging:
$\hat{\theta}_{n+k}=\theta_{n}-\sum_{i=1}^{k}w_{i}\cdot g_{n+i-1}$ (12)
This equation demonstrates that merging checkpoints with weights $\{c_{j}\}$ is equivalent to applying a synthetic decay schedule defined by weights $\{w_{i}\}$ to the gradients accumulated after the base checkpoint $\theta_{n}$, and there exists a mapping between the merging weights ${c_{j}}$ and the effective learning rates ${w_{i}}$
B.1 Proof of Theorem 3.1
We seek to find the unique checkpoint weights $\{c_{j}\}$ corresponding to a given desired sequence of gradient decay coefficients $\{w_{i}\}_{i=1}^{k}$. The relationship derived in the previous section is the starting point:
$w_{i}=\sum_{j=i}^{k}c_{j}$ (13)
We can solve for the checkpoint weights $\{c_{j}\}$ by starting from the last element and working backwards.
For $i=k$, the sum in Eq. 13 has only one term:
$w_{k}=\sum_{j=k}^{k}c_{j}=c_{k}$ (14)
This gives us the value of $c_{k}$ directly.
For any $i\in[1,k-1]$, we can write out the expressions for $w_{i}$ and $w_{i+1}$:
$w_{i}$ $=c_{i}+c_{i+1}+c_{i+2}+\dots+c_{k}$
$w_{i+1}$ $=\qquad c_{i+1}+c_{i+2}+\dots+c_{k}$
Subtracting the second equation from the first yields the expression for $c_{i}$:
$w_{i}-w_{i+1}=c_{i}$ (15)
Finally, for $c_{0}$, we use the constraint that the checkpoint weights must sum to one: $\sum_{j=0}^{k}c_{j}=1$.
$c_{0}$ $=1-\sum_{j=1}^{k}c_{j}$
$=1-\left(c_{1}+c_{2}+\dots+c_{k-1}+c_{k}\right)$
$=1-\left((w_{1}-w_{2})+(w_{2}-w_{3})+\dots+(w_{k-1}-w_{k})+w_{k}\right)$ (16)
The sum in the parentheses is a telescoping series which simplifies to $w_{1}$.
$c_{0}=1-w_{1}$ (17)
This completes the derivation of the unique formulas for $\{c_{j}\}$ as stated in the theorem.
For the checkpoint weights $\{c_{j}\}$ to be valid, they must be non-negative. This imposes conditions on the sequence $\{w_{i}\}$.
  • From Eq. 14, $c_{k}\geq 0$ implies $w_{k}\geq 0$.
  • From Eq. 15, $c_{j}\geq 0$ for $j\in[1,k-1]$ implies $w_{j}-w_{j+1}\geq 0$, which means $w_{j}\geq w_{j+1}$. This shows that the sequence $\{w_{i}\}$ must be monotonically non-increasing.
  • From Eq. 17, $c_{0}\geq 0$ implies $1-w_{1}\geq 0$, which means $w_{1}\leq 1$.
Combining these conditions, we arrive at the requirement that the gradient decay sequence must be bounded and monotonically non-increasing: $1\geq w_{1}\geq w_{2}\geq\dots\geq w_{k}\geq 0$. This ensures that a valid (non-negative) set of checkpoint weights $\{c_{j}\}$ can be derived.
Appendix C Additional Experiments
Our preliminary experiments into Warmup-Stable-Decay (WSD) learning rate schedules revealed a clear performance hierarchy among different decay curves, with the 1-sqrt strategy emerging as superior. Specifically, we conducted a controlled experiment initialized from a Ling-lite (Ling-Team et al., 2025) base checkpoint that was pre-trained on 7T tokens. We then annealed the model for an additional 100B tokens using a consistent, high-quality dataset, where we varied only the annealing decay function. The final learning rate for all experimental runs was set to zero. The results, depicted in Figure 7, confirm that the 1-sqrt decay outperforms other methods in benchmarks. Based on this evidence, we establish the WSD schedule with 1-sqrt decay as a strong baseline for all subsequent experiments. As this was a preliminary study, the starting checkpoint used here differs from that used in our main experiments in Section 4.
Figure 7: Comprehensive performance comparison between different decay strategy of WSM schedule.
Appendix D How Checkpoint Merging Influences Network Behavior: Stability & Generalizability
The architecture composed of sequentially connected transformer blocks has been widely adopted in mainstream LLMs. In such chain-like structures, variations in shallow-layer outputs typically propagate layer-wise. More stable and generalizable input-output patterns exhibit greater potential for enhancing downstream task performance. To investigate how checkpoint merging influences network behavior, we conduct the following analyses:
Stability
In numerical analysis, the condition number quantifies a function’s sensitivity to input perturbations and the resultant output errors. Remarkably, our analysis in Figure 8 reveals that the decay process induces sharp deterioration of condition numbers, whereas checkpoint merging demonstrates superior stability. This indicates that the checkpoint merging strategy not only improves performance but also preserves parametric stability.
Generalizability
When maintaining competitive performance, higher SVD entropy (singular values entropy) correlates with greater matrix effective rank, indicating greater information capacity in matrix operations. For continue pre-training and fine-tuning scenarios, higher SVD Entropy often means higher potential for model plasticity (Alter et al., 2000; Roy and Vetterli, 2007; Liu et al., 2025). Figure 9 shows the trend of SVD Entropy during training. We observe that decay is a rapid entropy reduction process, which continuously damages the potential for future continue training of the model. In contrast, the merged models still maintain a higher generalizability, manifested as a higher SVD Entropy.
Figure 8: Condition number of weight matrices across different training tokens.
Figure 9: The entropy of singular values for each weight matrix in the feed-forward network (FFN) layers.
Appendix E Detailed Evaluation Results
We provide detailed evaluation results to compare our methods with WSD. Figure 10 shows a detailed comparison on each dataset across the various categories from the main experiments in Section 4.2, where WSM achieves an advantage on the vast majority of datasets. We select the checkpoint with the highest average score for each method, including WSD and the three WSM merging algorithms, and list them in Table 7. Table 8 and 9 shows the performance comparison of checkpoints generated by the WSM and WSD schedules after supervised fine-tuning (SFT).
Figure 10: Detailed performance comparison between the standard WSD schedule (via learning rate decay) and our WSM schedule (via checkpoint merging) under different merging durations.
Table 7: Detailed performance comparison of base models trained using WSM (with three distinct merging algorithms) versus WSD scheduling approaches.
Metric WSD WSM
EMA mean 1-sqrt
General Knowledge & Reasoning ARC-c 88.8 87.8 88.1 89.8
ARC-e 95.4 93.7 94.0 94.5
AGIEval 53.2 53.8 54.4 54.2
OpenBookQA_fact 89.2 89.2 92.2 91.8
BBH 69.0 70.0 71.2 71.0
BBH-zh 64.5 65.3 67.3 66.5
WorldSense 50.4 50.6 51.9 52.1
PIQA 78.6 78.2 78.5 78.5
HellaSwag 69.6 68.2 69.9 69.7
KOR-Bench 31.9 33.7 34.8 34.6
Language Understanding race-middle 85.7 88.7 88.8 87.9
race-high 81.9 82.2 84.2 83.9
SQuAD2.0 82.9 81.6 81.5 81.3
TriviaQA_5shot 58.8 57.5 59.8 59.7
NQ_5shot 29.8 29.1 30.4 30.3
WinoGrande 67.6 66.7 67.3 66.6
Professional Knowledge MMLU 70.7 71.2 72.2 72.1
GPQA 24.8 30.8 33.3 30.8
CMMLU 79.6 78.3 79.2 79.3
MMLU-Pro 43.3 43.8 45.2 44.7
SuperGPQA 23.8 24.3 25.5 25.7
C-Eval 78.5 78.3 80.9 79.9
Math gsm8k 83.6 82.0 82.6 82.3
MATH 48.3 50.8 50.0 49.6
gaokao 38.5 47.3 39.6 48.4
gaokao2023en 54.0 49.1 53.5 53.0
gsm_plus 71.5 69.3 71.4 70.9
mgsm_zh 72.0 73.2 73.2 72.8
CMATH 88.6 88.4 89.3 89.2
MathBench 57.6 61.8 60.3 61.6
minerva_math 16.9 19.9 20.6 19.1
college_math 38.9 37.8 38.3 39.0
cn_middle_school_24 62.4 67.3 68.3 70.3
Code HumanEval 79.9 80.5 81.7 81.1
LiveCodeBench 21.2 20.3 23.2 22.7
HumanEval_plus 75.0 75.6 77.4 77.4
HumanEval_fim 73.8 69.3 73.1 73.9
HumanEval_cn 74.4 77.4 75.0 75.6
MBPP 67.2 65.0 66.8 66.8
MBPP_plus 72.0 70.6 70.4 70.6
CruxEval 55.6 54.8 57.0 57.4
Table 8: Detailed performance comparison of checkpoints generated by the WSM and WSD schedule after supervised fine-tuning (SFT). Both base checkpoints are fine-tuned under identical settings for 5 epochs. Results are reported based on the epoch with the highest average benchmark score.
Metric WSD WSM
Knowledge Basic Knowledge ARC-c 90.85 89.49
BoolQ 84.80 85.38
GaokaoBench 75.70 79.95
AGIEval 61.87 65.22
NQ 25.4 26.43
TriviaQA 53.93 55.52
Average 65.42 67.00
Professional Knowledge C-Eval 76.37 77.87
CMMLU 76.13 76.78
MMLU 72.76 74.59
MMLU-Pro 46.09 49.67
GPQA 29.55 33.4
SuperGQPA 26.50 26.43
Average 54.57 56.46
Code Code Completion HumanEval 85.90 86.20
HumanEval_plus 81.10 80.95
HumanEval_cn 78.20 76.07
CruxEval 59.69 58.88
Multiple 66.14 64.75
HumanEvalFix 63.01 62.50
Average 72.34 71.56
Code Generation MBPP 81.03 81.97
MBPP_plus 73.02 71.43
LiveCodeBench 31.31 33.31
BigCodeBench 33.77 33.16
CodeForces 19.60 18.07
Bird-SQL 25.95 28.10
Average 44.11 44.34
Math Elementary Mathematics CMATH 93.08 93.99
gsm8k 87.34 87.79
cn_middle_school_24 72.28 73.27
mgsm_zh 76.80 78.25
gsm_plus 77.28 77.93
Average 81.36 82.25
Intermediate Mathematics MATH 74.24 77.08
MathBench 74.72 75.74
college math 43.57 43.52
gaokao 64.17 60.77
minerva math 55.74 57.31
gaokao2023en 63.12 65.19
MATH500 74.75 77.05
Average 64.33 65.24
Advanced Mathematics OlympiadBench 41.89 41.93
AIME2024 11.04 11.67
AIME2025 11.46 12.71
Average 21.46 22.10
Table 9: Detailed performance comparison of checkpoints generated by the WSM and WSD schedule after supervised fine-tuning (SFT). Both base checkpoints are fine-tuned under identical settings for 5 epochs. Results are reported based on the epoch with the highest average benchmark score.
Metric WSD WSM
Language Language Understanding C3 84.38 87.18
WSC 69.23 77.88
race-high 82.93 84.65
race-middle 87.95 89.42
Average 81.12 84.78
Reasoning Complex Reasoning bbh 72.87 73.97
drop 76.41 78.66
hellaswag 68.11 70.93
ocnli 51.32 52.75
piqa 81.50 82.92
ProntoQA 37.00 38.00
Multi-LogiEval 0.27 56.68
MuSR 48.53 51.33
korbench 37.52 37.52
bbh-zh 69.38 71.47
Average 63.21 64.94
Agent Tool-use teval_v2_en 84.4 86.16
teval_v2_zh 83.49 84.72
BFCL_AST 79.35 78.30
BFCL-Live 68.33 71.39
NEXUS 30.51 29.21
Average 69.22 69.96
Instruction Following IFEval 71.9 75.71
alignbench 59.10 59.80
Average 65.5 67.75